### v1

# Interpretability Needs a New Paradigm

Abstract
Interpretability is the study of explaining models in understandable terms to humans. At present, interpretability is divided into two paradigms: the intrinsic paradigm, which believes that only models designed to be explained can be explained, and the post-hoc paradigm, which believes that black-box models can be explained. At the core of this debate is how each paradigm ensures its explanations are faithful, i.e., true to the model’s behavior. This is important, as false but convincing explanations lead to unsupported confidence in artificial intelligence (AI), which can be dangerous. This paper’s position is that we should think about new paradigms while staying vigilant regarding faithfulness. First, by examining the history of paradigms in science, we see that paradigms are constantly evolving. Then, by examining the current paradigms, we can understand their underlying beliefs, the value they bring, and their limitations. Finally, this paper presents 3 emerging paradigms for interpretability. The first paradigm designs models such that faithfulness can be easily measured. Another optimizes models such that explanations become faithful. The last paradigm proposes to develop models that produce both a prediction and an explanation.
Interpretability, Explanations, Transparency, Paradigms, Post-hoc, Intrinsic, Ethics, Future work, Faithfulness measurable models, Self-explanations, Self-explaining models
1 Introduction
There was a time in physics, in the late 17th century, when Isaac Newton insisted that light is a particle and Christiaan Huygens insisted that light is a wave (Huygens, 1690). These ideas were seemingly irreconcilable at the time. Of course, now we have a much better theory, and we understand that light can be seen as both a wave and a particle. [^1] Known as the wave-particle duality concept (Messiah, 1966).
In 1874, Georg Cantor proposed set theory and showed there exists at least two kinds of infinity. This divided the mathematical field. The Intuitionists, who named Cantor’s theory nonsense, thought that math was a pure creation of the mind and that these infinities weren’t real. Henri Poincaré said: “Later generations will regard Mengenlehre (set theory) as a disease from which one has recovered” (Gray, 1991). Leopold Kronecker called Cantor a “scientific charlatan” and “corruptor of the youth” (Dauben, 1977).
The other group, the Formalists, thought that by using Cantor’s set theory, all math could be proven from this fundamental foundation. David Hilbert said: “No one shall expel us from the paradise that Candor has created” and “In opposition to the foolish Ignoramus (we will not know; i.e., intuitionists), our slogan shall be: We must know – we will know” (Smith, 2014).
Today, we know infinities are important concepts; thus, the Intuitionists were wrong. However, Kurt Gödel showed that the Formalists were also wrong. Unfortunately, there exist true statements which can never be proven (Gödel, 1931, Gödel’s incompleteness theorem).
These are just two examples in science and mathematics where there have been strong debates and beliefs due to conflicting paradigms. Science historian Thomas Kuhn defines a scientific paradigm as: “universally recognized scientific achievements that, for a time, provide model problems and solutions to a community of practitioners” (Kuhn, 1996).
Time and time again, when there are conflicting paradigms, it is only “for a time”. Eventually, we find neither paradigm is true, or both paradigms are true (under a more nuanced understanding). In retrospect, it is more constructive to develop an understanding as to which paradigms may be right under what conditions, as opposed to an all-or-nothing approach of arguing about a singular right paradigm. Alternatively, we could come up with a new paradigm, a new school of thought, a new direction; which replaces or bridges the old way of thinking.
In this paper, we re-examine the current direction and paradigms of interpretability and invite the reader to consider whether it is time for a new paradigm.
1.1 Interpretability and faithfulness
Interpretability is the ability to explain a model in understandable terms to humans (Doshi-Velez & Kim, 2017). Model explanations have particularly become important for AI safety, as machine learning is increasingly being used by the industry and affects the lives of most humans. This and additional motivations are elaborated on in Section 2.
Within interpretability, there currently exist two paradigms, called post-hoc and intrinsic (Lipton, 2018). Section 3 properly describe their stance. Put briefly, the intrinsic paradigm believes that only models designed to be explained can be explained (Rudin, 2019);. In contrast, the post-hoc paradigms believe this constraint is unnecessary and too restrictive to achieve competitive performance (Madsen et al., 2022b).
The position in this paper is that neither paradigm has been fruitful because their underlying beliefs are problematic and we should therefore look for new directions. Section 4 contains the primary support for this position. To prove that other paradigms can exist, Section 5 then presents three emerging paradigms for interpretability and discusses how they might overcome past challenges, their beliefs, drawbacks, and future directions. Importantly, while this demonstrates that developing new paradigms is possible, Section 5 should not be considered a final list of directions.
At the core of this discussion is how each paradigm approaches faithfulness. A faithful explanation means the explanation accurately reflects the model’s logic, and ensuring and validating this often presents a major challenge because the model’s logic is inaccessible to humans (Jacovi & Goldberg, 2020). Faithfulness is particularly important, as false but convincing explanations can lead to unsupported confidence in models, increasing the risk of AI.
New paradigms, such as those described in Section 4, bring new perspectives regarding how to achieve faithfulness. This creates a new opportunity to do interpretability research centered around ensuring faithfulness. However, it also creates a new risk as we may take faithfulness for granted once again, as has been the case with both the intrinsic (Jacovi & Goldberg, 2020) and post-hoc paradigms (Madsen et al., 2022b). To prevent this, this paper also takes the position that we should be vigilant about faithfulness when it comes to these new paradigms to prevent repeating past mistakes.
2 Why interpretability is needed
Before discussing the current paradigms and their shortcomings, it’s necessary to first consider if interpretability is needed at all. Many ethical motivations for interpretability are also served by bias and fairness metrics, so if the current paradigms of interpretability do not work (as we argue in Section 4), perhaps we should drop the idea of interpretability completely. If the models can be made accurate, unbiased, and fair enough, do we need to explain the models? In this section, we will argue that interpretability is required, by examining the limitations of bias and fairness metrics and the scientific motivations for interpretability.
2.1 Limitations of bias and fairness metrics
There is no doubt that bias and fairness metrics present a vital role in validating models’ behavior. However, a shared limitation is that they always measure known attributes (Barocas et al., 2019). For example, gender-bias metrics use gender attributes. This presents two challenges. Can we procure such attributes (known as protected attributes)? How do we prevent unanticipated biases?
2.1.1 Protected attribute procurement
Attributes like gender, race, age, disability, etc., are under U.S. law known as “protected attributes” (Xiang & Raji, 2019), and collecting and using these attributes is heavily regulated in most of the world. Andrus et al. (2021) writes, “In many situations, however, information about demographics can be extremely difficult for practitioners to even procure.”. Therefore, systematically measuring bias and fairness is not always practical (Andrus et al., 2021).
On the other hand, explanations often don’t depend on knowing these protected attributes in advance and can provide a more qualitative analysis. For example, suppose an explanation tells us that the word “Woman” from “Member of Woman’s Chess Club” in a resume is important for making a hiring recommendation. In that case, there is a potential harmful bias (Kodiyan, 2019). Therefore, explanations can serve a similar practical purpose to a fairness or bias metric without performing systematical correlations.
2.1.2 Unknown attribute bias
Although protected attributes are important to consider and are often legally protected, many more relevant attributes are involved in ensuring a fair and unbiased system. Unfortunately, it is impossible to consider every possible bias in advance. As an alternative, interpretability offers a more qualitative and explorative validation.
Continuing the example with resumes and automated hiring recommendations, during investigations by Fuller et al. (2021), the authors found that a hospital only accepted candidates with computer programming experience when they needed workers to enter patient data into a computer. Another example was a clerk position where applicants were rejected if they did not mention floor-buffing (i.e., a cleaning method for floors) (Fuller, 2021).
These examples present cases of systematic unintended bias. However, they do not relate to any protected attributes, and they are so specific they can only be discovered through qualitative explanations and investigations. That said, systematic fairness/bias metrics can quantify the damage once potential biases are identified using interpretability. Afterwards, those metrics can be integrated into a quality assessment system to prevent future harm.
2.2 Interpretability for scientific discovery and understanding
Interpretability is not only used for ethics and adjacent purposes, where bias and fairness metrics have an important role. Interpretability is also used for scientific discovery and learning about what makes models work.
2.2.1 Scientific Discovery
An example of scientific discovery is interpretability in drug discovery (Preuer et al., 2019; Jiménez-Luna et al., 2020; Dara et al., 2022). A common approach is to use importance measures to identify regions in genomic sequences responsible for a particular behavior, such as producing a protein. While these explanations do not guarantee that such connections exist in reality, they can provide important initial hypotheses for scientists enabling them to make more informed choices about the direction of their research.
2.2.2 Model understanding
An emerging field of interpretability is mechanistic interpretability, which identifies parts of a neural network that have a particular responsibility (Cammarata et al., 2020). For example, identifying a collection of neurons responsible for copying content in a generative language model, etc. (Elhage et al., 2021). Such insights may not be directly relevant to downstream tasks, but they help us understand current model limitations and can lead to better model design.
3 The current paradigms of interpretability
Table 1: Comparison of the definitions and underlying beliefs of the intrinsic and post-hoc paradigms. The beliefs relate to a) requirements for a faithful explanation and b) model capabilities. It should be apparent that these two views are seemingly incompatible.
[…]
This paper uses a common definition of interpretability, “the ability to explain or to present in understandable terms to a human” by (Doshi-Velez & Kim, 2017). However, even this definition of interpretability is not agreed upon.
Lipton says, “the term interpretability holds no agreed upon meaning, and yet machine learning conferences frequently publish papers which wield the term in a quasi-mathematical way” (Lipton, 2018). In 2017, a UK Government House of Lords review of AI noted after substantial expert evidence that “the terminology used by our witnesses varied widely. Many used the term transparency, while others used interpretability or explainability, sometimes interchangeably” (House of Lords, 2017, 91).
For this reason, there are also no clearly agreed-upon definitions of the current paradigms of interpretability (Carvalho et al., 2019; Flora et al., 2022). As such, this section defines the intrinsic and post-hoc paradigms, as well as describe their underlying beliefs, which are summarized in Table 1.
3.1 Definitions
Jacovi & Goldberg (2020) write: “A distinction is often made between two methods of interpretability: (1) interpreting existing models via post-hoc techniques; and (2) designing inherently interpretable models. (Rudin, 2019)”. Based on this and other sources (Madsen et al., 2022b; Arya et al., 2019; Carvalho et al., 2019; Murdoch et al., 2019), this paper refers to these two ideas as 1) the intrinsic paradigm and 2) the post-hoc paradigm.
3.1.1 The intrinsic paradigm
The intrinsic paradigm works on creating so-called inherently interpretable models. These models are architecturally constrained, such that the explanation emerges from the architecture itself.
Classical examples are decision trees or linear regression. In the field of neural networks some examples are: 1) “Old-school” attention (Bahdanau et al., 2015; Jain & Wallace, 2019), where attention points to which input tokens are important. 2) Neural Modular Networks (Andreas et al., 2016; Gupta et al., 2020; Fashandi, 2023), which produce a prediction via a sequence of sub-models, each with known behavior. 3) Prototypical Networks (Bien & Tibshirani, 2009; Kim et al., 2014; Chen et al., 2019), which predicts by finding similar training observations.
3.1.2 The post-hoc paradigm
Post-hoc explanations are computed after the model has been trained. They are developed independently of the model’s architecture and how it was trained. However, there are often some simple criteria, like “the model should be differentiable”, “the training dataset is known”, or “inputs are represented as tokens” (Madsen et al., 2022b). Although general applicability is technically not a requirement, if a method is so specific that it only works on one specific model, it’s likely an intrinsic explanation.
As an example, a common post-hoc explanation is gradient-based importance measures. Importance measures explain which input features (words, pixels, etc.) are important for making a prediction. This is archived by differentiating the prediction with respect to the input. The idea is that if a small change in input causes a big change in the output, then that input is important (Baehrens et al., 2010; Seo et al., 2018; Karpathy et al., 2015).
3.2 Beliefs
As with all paradigms, there are fundamental underlying beliefs, which are why the paradigm’s followers partake in their paradigm of choice. At the core of these beliefs are two central questions. When are explanations faithful and what are the requirements for faithfulness? And, how do these requirements affect the model’s general performance capabilities?
3.2.1 When are explanations faithful?
The intrinsic paradigm believes that: only models designed to be explained, can be explained, which their inherently interpretable models try to satisfy. Therefore, they argue that using black-box models is too risky, as these models can never be faithfully explained (Rudin, 2019).
However, although their models are designed to be intrinsically explainable, this claim and their faithfulness should still be questioned (Jacovi & Goldberg, 2020), as many inherently interpretable model ideas are later revealed not to provide faithful explanations. For example, attention-based explanations have received notable criticism for not being faithful (Jain & Wallace, 2019; Serrano & Smith, 2019; Vashishth et al., 2019; Meister et al., 2021; Madsen et al., 2022a). This is discussed more in Section 4.1.
The post-hoc explanation paradigm takes a less strict stance and believes that even models that were not designed to be explained (i.e., black-box models) can still be explained. However, as this paradigm has no control over the model, achieving faithful explanations is very challenging; this is discussed more in Section 4.2.
In conclusion, the intrinsic paradigm considers explanations to be part of the model design, and post-hoc explanations are always applied after the model design. Madsen et al. (2022b) frame intrinsic as proactive and post-hoc as retroactive. Hence, the two schools of thought are incompatible frameworks, and they can philosophically be considered as paradigms (Kuhn, 1996).
3.2.2 What is the effect on the model’s general performance capabilities?
It would seem that intrinsic explanation is the obvious choice. If we can control the model such that the faithfulness of explanations can be guaranteed, why consider post-hoc explanation?
The commonly mentioned idea is that the post-hoc paradigm believes that by constraining the models in the manners that the intrinsic paradigm requires, there is a trade-off in performance (DARPA, 2016). However, this trade-off does not have to be the case in practice (Rudin, 2019, section 2).
A more accurate take, which is rarely explicitly discussed, is that the common industry prefers off-the-shelf general-purpose models and only later thinks about interpretability (Bhatt et al., 2019). Additionally, most research only considers predictive performance, not interpretability. Therefore, intrinsic researchers are always catching up to black-box models. From the post-hoc perspective, it would make more sense to work on generally applicable interpretability methods for both off-the-shelf and future black-box models.
From the intrinsic perspective, while the industry might prefer off-the-shelf models now, they shouldn’t. Not validating models through intrinsic explanations can have serious consequences (Rudin, 2019) and eventually damage their business. Additionally, with increasing legal requirements to provide explanations, the industry may have to use inherently explainable models (Goodman & Flaxman, 2017).
For these reasons, the intrinsic paradigm believes we should not let the industry’s needs dictate our research direction, as their goals may be too short-sighted. In the long run, intrinsic models may be the only reasonable option.
In conclusion, the post-hoc paradigm has good intentions with providing general explanations for general-purpose models. However, from the intrinsic paradigm perspective, those good intentions are meaningless if it is fundamentally impossible to provide guaranteed faithful explanations without an inherently interpretable model.
4 Why interpretability needs a new paradigm
It tends to be the case that when there are multiple paradigms, it is because neither of the paradigms fits the needs. However, for the case of the post-hoc and intrinsic paradigms, it could be argued that they serve different needs. For example, intrinsic explanations should be preferred for critical applications (Rudin, 2019), and post-hoc explanations could be used for verifiable situations, such as drug discovery, where the hypothesis generated by the explanations is verified using physical experiments.
4.1 The case against the intrinsic paradigm
The industry primarily uses post-hoc explanations, including for high-stakes applications such as insurance risk assessment and financial loan assessment (Bhatt et al., 2019; Krishna et al., 2022). This is because such industries usually do not have the in-house expertise to develop custom inherently interpretable models. They must rely on basic inherently interpretable models, like decision trees, which are not competitive, or use more advanced off-the-shelf neural black-box models, like pre-trained language models, which will be competitive. In practice, the industry is thus not in a position to choose inherently interpretable models.
Another challenge with the intrinsic paradigm is that its models are often not completely interpretable because only a part of the model is aritectually constrained to be interpretable. The rest of the neural network, still use black-box components (e.g. Dense layer, Recurrent layer, etc.) which are not interpretable. As such, the intrinsic promise should not be taken at face value (Jacovi & Goldberg, 2020).
An example of this, is classic attention-based models (Bahdanau et al., 2015; Jain & Wallace, 2019). Attention itself is interpretable, as it’s a weighted sum, and explains the importance of each intermediate representation. However, attention is often used as token-importance. This does not work, as the intermediate representations are produced by a black-box recurrent neural network (e.g., LSTM Hochreiter & Schmidhuber 1997) which can mix or move the relationship between tokens and the intermediate representations. Therefore, the attention scores do not necessary represent token-importance (Bastings & Filippova, 2020).
Likewise, Neural Modular Networks produce an executable problem composed of sub-networks, such as find-max-num(filter(find())), which is interpretable (Fashandi, 2023; Andreas et al., 2016; Gupta et al., 2020). However, each sub-networks (find-max-num, filter, find) is itself a black-box model with little guarantee that it operates as intended (Amer & Maul, 2019; Subramanian et al., 2020; Lyu et al., 2022).
Overall, there are few success stories with intrinsic explanations. They are either not performance-wise competitive, general-purpose enough for the industry (Bhatt et al., 2019), or their intrinsic claims are unsupported (Jacovi & Goldberg, 2020).
4.2 The case against the post-hoc paradigm
Although post-hoc explanations directly address the interpretability challenge of black-box components and models, and could therefore provide more complete explanations, there are also few success stories with post-hoc, where post-hoc explanations are consistently faithful.
Most notable is perhaps post-hoc importance measure (IM) explanations, where the explanation indicates which input features are the most important for making a prediction. The pursuit of such explanations have produced countless papers (Binder et al., 2016; Ribeiro et al., 2016; Li et al., 2016; Shrikumar et al., 2017; Smilkov et al., 2017; Sundararajan et al., 2017; Ahern et al., 2019; Thorne et al., 2019; ElShawi et al., 2019; Sangroya et al., 2020).
However, repeatedly, the faithfulness of these IM explanations is criticized (Adebayo et al., 2018, 2021; Kindermans et al., 2019; Hooker et al., 2019; Slack et al., 2020; Yeh et al., 2019). For example, there is great disagreement between alleged faithful IMs, which is hard to reconcile (Jain & Wallace, 2019; Krishna et al., 2022). Other works show their faithfulness is both task and model-dependent and thus don’t provide the generality that the post-hoc paradigm desires (Bastings et al., 2022; Madsen et al., 2022a). Finally, theoretical works suggest that IMs are subject to a no free lunch thorem (Han et al., 2022), or it may be impossible to provide faithful post-hoc IMs (Bilodeau et al., 2024).
Similar to the work of IM, is visualization of neurons in computer vision, which shows that neurons represent high-level concepts, such as nose or dog. This is done by visualizing convolutional weights or the input image that maximizes a neuron’s activation (Olah et al., 2017; Nguyen et al., 2016; Yosinski et al., 2015), which provides very convincing evidence. However, it has been shown empirically, theoretically, and through human-computer-interaction (HCI) studies that these visualizations do not provide useful explanations regarding the neurons’ responsibility (Geirhos et al., 2023; Borowski et al., 2021; Zimmermann et al., 2021) [^2] Neural networks likely do encode high-level concepts, but these visualizations are not useful for identifying the responsibility of specific neurons..
Another notable example is probing explanations, where models are verified by relating the model’s behavior or intermediate representation to, for example, linguistic properties (part-of-speech, etc.) (Belinkov & Glass, 2019; Belinkov et al., 2020). This idea has produced an entire subfield called BERTology (Rogers et al., 2020). BERTology in particular has attained substantial attention (Coenen et al., 2019; Clark et al., 2019; Rogers et al., 2020; Clouatre et al., 2022; McCoy et al., 2019; Conneau et al., 2018; Tenney et al., 2019), with most of the works finding that neural networks can learn linguistic properties indirectly.
Unfortunately, like post-hoc importance measures, there are many reasons to be highly skeptical (Belinkov, 2021). For example, using an untrained model or a randomized dataset shows an equally high correlation with linguistic properties, compared with training a regular model (Zhang & Bowman, 2018; Hewitt & Liang, 2019). These discoveries have put the entire methodology into question, although there is work trying to adapt to these new critics (Voita & Titov, 2020).
4.3 Conclusion
Post-hoc importance measures and probing explanations are just two cases where post-hoc shows initial promise through countless papers, only to be debunked repeatedly. The trend is oscillating between proposing new explanation methods and debunking them. Of course, it’s impossible to prove that there will never be a great post-hoc method. However, the lack of guarantees also makes it impossible to know when a faithful post-hoc method is proposed. Similarly, intrinsic explanations also receive criticism after a while, as has been the case with attention and Neural Modular Networks.
Although both the intrinsic and post-hoc paradigms have significant issues, parts of their underlying beliefs have merit. The intrinsic paradigm believes that we can’t expect models that were not designed to be explained to be explained, while post-hoc believes black-box models tend to be more general purpose and while providing high predictive performance. These beliefs have merit, and it’s worth considering how to incorporate their spirit into new paradigms.
5 Possible new paradigms
Table 2: Comparison of the definitions and underlying beliefs of the new paradigms. The beliefs relate to a) explanation requirements and b) model capabilities. These new paradigms can be compared with the old paradigms in Table 1.
[…]
Given the shortcomings of the intrinsic and post-hoc paradigms, we are now in a situation where it’s meaningful to look beyond these current paradigms. To prove that developing alternative paradigms is possible and to highlight some new directions of research, this section discusses new directions for interpretability, frames them as paradigms, and summarizes them in Table 2. However, these new paradigms are not the only directions one could take for interpretability, and in general, we will encourage researchers to consider alternatives that have not yet been explored.
5.1 Inherently Faithfulness Measurable Models
The faithfulness of an explanation can be extremely difficult to measure because the true explanation is inaccessible to us (Jacovi & Goldberg, 2020). Therefore, faithfulness metrics often have to make use of a proxy measure.
For example, the erasure metric for importance measures says: if information (pixels, tokens, etc.) is truly important, then when removing it the prediction should change significantly (Samek et al., 2017). However, this has various limitations, because removing information causes out-of-distribution issues, which other works have solved by repeatedly retraining the model on datasets with allegedly important information removed (Hooker et al., 2019). Unfortunately, retraining is computationally expensive, and measuring on a retrained model means faithfulness is no longer measured on the original model. Other faithfulness metrics have similar limitations (Madsen et al., 2024b).
Therefore, rather than considering a faithfulness metric as something that exists externally of the model, the idea with Inherently Faithfulness Measurable Models (FMMs) is a model that inherently provides the means to cheaply and accurately measure faithfulness (Madsen et al., 2024b). Then, because the faithfulness measure is cheap and accurate, it’s possible to optimize an explanation to maximize this faithfulness metric (Zhou & Shah, 2023).
This is different from how inherently explainable models in the intrinsic paradigm which are directly designed to be explained because their architectural constraints produce the explanation (Jain & Wallace, 2019; Rudin, 2019). The central idea in FFMs is the model does not produce an explanation directly but enables a faithfulness metric, thus making FFMs indirectly designed to be explainable (Madsen et al., 2024b).
In essence, Madsen et al. (2024b) reformulates the interpretability problem from ‘inherently explainable” to “inherently measurable” and show that this reformulation is much easier to design for, as it does not require architectural constraints. The concrete demonstration that Madsen et al. (2024b) provide, uses a RoBERTa model (Liu et al., 2019) where the loss function is modified such that removing tokens becomes in-distribution, thereby enabling the erasure metric (Samek et al., 2017) to be applied directly without retraining. This achieves state-of-the-art faithfulness metrics without affecting performance or inference time.
Because the work done by Madsen et al. (2024b) only applies to masked language models. Thus, the immediate next step is to extend these ideas to causal language models which may be a challenging task as their idea depends on mask-tokens. Additionally, new work may explore more ways of optimizing explanations given a cheap faithfulness metric (Zhou & Shah, 2023).
Beyond these immediate applications, the principle of enabling models to be faithfulness measurable can be applied to other explanations. For example, concepts (Kim et al., 2018) and influential examples (Yeh et al., 2018; Pruthi et al., 2020; Koh & Liang, 2017) may be good candidates for this. Both the general idea and the specific approach by Madsen et al. (2024b), can also be applied to other domains such as computer vision.
To summarize, while previous interpretability works have explained first and then measure faithfulness later, this paradigm does the opposite; it designs for faithfulness measures first and then explains. This reformulation of the interpretability problem is easier to design for, thus satifing both the black-box and designed-to-be-explained desirables.
5.2 Learn to faithfully explain
Rather than optimizing the predictive model to be faithfulness measurable, a more long-standing idea in interpretability has been to optimize the predictive model such that an explanation method applied to that model produces faithful explanations (Yoon et al., 2019; Chen et al., 2018). This also does not require the architectural constraints that the intrinsic paradigm applies, as the explanation comes from an external explanation method, not the architectural design.
The explanation method can be similar or even identical to those from the post-hoc paradigm. However, because the model is optimized to enable these explanations to be faithful, it’s not post-hoc, and there are more reasons to think that the explanations should be faithful.
Early work on this used an explanation model, which will produce an explanation of the prediction model. The explanation and prediction models are then trained jointly to produce faithful explanations and high-accuracy predictions (Yoon et al., 2019; Chen et al., 2018). This direction has been called joint amortized explanation methods (JAMs). However, Jethani et al. (2021) point out that the explanation model often learns to encode the prediction, which means the explanation model becomes part of the black-box problem rather than the solution.
Jethani et al. (2021) solve this by using a disjoint setup, where the explanation model can’t encode the prediction, a setup that following works have adapted (Jethani et al., 2022; Covert et al., 2022). However, this still presents a problem as it assumes the explanation model is 100% robust. For example, an out-of-distribution observation could break the explanation model and produce unfaithful explanations. This is particularly relevant, as it is often the cases where models break that are desirable to explain.
Instead, a better direction is likely to use a simple explanation algorithm that is easy to understand rather than an explanation model. Bhalla et al. (2023) do this by jointly training a predictive model with an explanation algorithm, where explanations indicate unnecessary features and the model learns to support sparse features.
Regardless of the specific approach used to produce the explanation, the optimization criteria used to make the predictive model converge, such that the explanations are faithful, is either a relaxation of a faithfulness metric or an underlying criterion necessary to produce faithful explanations. For example, masking inputs should be in-distribution (Covert et al., 2022; Bhalla et al., 2023). Therefore, this field needs to work on how to ensure and validate that the explanations are actually faithful and that the model properties that make explanations faithful also hold for out-of-distribution data.
Another direction in this paradigm is the idea that regularization can produce more faithful explanations. For example, a more locally-linear model may be more suitable for a gradient-based explanation (Ross & Doshi-Velez, 2018; Srinivas et al., 2022; Geirhos et al., 2023), or training a computer vision model with added input-noise reduces noise in importance measure explanations because the model becomes less sensitive to one-pixel variations (Bansal et al., 2020). As of now, these observations are mostly hypothetical or anecdotal, and we advocate for a systematic study from an interpretability perspective.
5.3 Self-explaining models
Rather than using external algorithms or models to produce explanations, Elton (2020) proposes that models should explain themselves, meaning they become self-explaining. Because there are now two black-box modalities (the self-explaining and the normal prediction modality), this can be quite dangerous, and there must be solid evidence that the explanations are faithful for this approach to be valid.
Despite the immediate dangers with a self-explaining model, the best tool to explain a black-box model may be the model itself. In particular, unlike the JAM approach (mentioned in Section 5.2), a self-explaining model has full access to its own internal weights and activations. The explanations can also be much more informative. This particularly holds for instruction-tuned large-language models (LLMs) (OpenAI, 2023; Jiang et al., 2023; Meta, 2023), which are allegedly able to explain themselves in great detail and very convincingly (Chen et al., 2023; Agarwal et al., 2024).
When LLMs explain themselves, this is referred to as self-explanations. Importantly, self-explanations must relate to the model’s reasoning logic, not just the world or abstract concepts (e.g., mathematics), which is how most Chain-of-Thought (CoT) explanations are evaluated (Chu et al., 2023). As humans don’t know the model’s logic, evaluating self-explanations is difficult (Huang et al., 2023).
Chen et al. (2023) show although self-explanations are very convincing to humans (called plausibility), they do not help humans to guess model predictions of new observations (called simulatability) and, therefore, can be dangerous. They argue that “naively optimizing human approvals (e.g., RLHF)” does not improve self-explanations’ simulatability.
Simulatability is important for measuring the usefulness of explanations. However, it is difficult to ensure that the human test subjects make their guesses based on the model’s self-explanation and not their world knowledge. Therefore, measuring faithfulness is equally important, as this evaluates explanations independently of humans.
To measure faithfulness in LLMs, one can use self-consistency checks (Madsen et al., 2024a; Huang et al., 2023; Lanham et al., 2023). For example, Lanham et al. (2023) considers CoT explanations for mathematical reasoning, edits facts in the CoT (e.g., from “2+2=4” to “2+3=5”), and then checks that the prediction changes based on these edits. Another example is Madsen et al. (2024a), who evaluate counterfactual and importance measure explanations by asking a model to change the input such it gives a different prediction; this can be verified by re-evaluating with the counterfactual input. Both Lanham et al. (2023) and Madsen et al. (2024a) find that the faithfulness of LLMs’ self-explanations is model and task-dependent.
Unfortunately, self-consistency checks are impossible for most explanations (Madsen et al., 2024a; Lanham et al., 2023). An alternative may be evaluating the model’s ability to model itself (called self-modeling). The hypothesis is that self-modeling is a necessary condition for general faithful self-explanations (Madsen et al., 2024a).
Anthropic Team (2022) evaluate self-modeling capabilities (not in a self-explanation context) by asking the LLM if it knows the answer to a question and comparing it with simply asking the LLM to answer the question. They find that LLMs cannot perform this task without additional fine-tuning. Even with fine-tuning, the LLM cannot generalize to other domains. As such, developing general self-modeling capabilities could be a first step towards general self-explanation. Such work should also show that self-modeling capabilities make concrete explanations faithful too, using self-consistency checks (Madsen et al., 2024a).
5.4 Comparison
All three mentioned paradigms work with what-would-be black-box models. However, their idea is to optimize these models to guarantee a faithful explanation. They optimize for either faithfulness measurability, explanation faithfulness, or self-model/self-explanation capabilities.
In conclusion, they are different from the post-hoc paradigm because interpretability is considered during optimization, and they are different from the intrinsic paradigm because they are not architecturally constrained. They are also mutually distinct, each taking a different perspective on achieving faithful interpretability via optimization.
At first, this sounds like a natural and simple idea. However, recall that we can’t know what a faithful explanation is or what it looks like, as the model’s internal logic is inaccessible to us. Achieving interpretability via optimization is thus extremely challenging and involves proxies, assumptions, etc., which is how they differ.
6 Limitations
As with any other position paper, this paper presents one position: that we must seek out new paradigms of interpretability. However, one could make other valid positions or directions about the future of interpretability.
In particular, this position paper primarily focuses on faithfulness. However, as interpretability is about how we can explain models to humans in understandable terms, one must also consider how well-understandable an explanation is. This concern is known as human-groundedness (Doshi-Velez & Kim, 2017), simulatability (Lipton, 2018), or comprehensibility (Robnik-Šikonja & Bohanec, 2018).
This concern is orthogonal to the faithfulness concern and is therefore not discussed in this paper. However, readers are encouraged to study recent works like Schut et al. (2023); Kim (2022), which proposes the new idea that it is not enough to frame explanations in terms that humans already understand. We should also develop new language and mental abstractions for humans to understand machines.
7 Conclusion
Although some evidence exists for the new paradigms that were presented in Section 5, these are, first and foremost, just ideas. It’s only in retrospect that we can truly know if one paradigm results in meaningful progress in the field. It is also entirely possible that neither of these ideas is what moves the interpretability field forward.
For these reasons, the core position of this paper is that we should develop new directions and paradigms in interpretability, instead of focusing on the existing post-hoc and intinsic paradigms, which are currently dominating.
That being said, we must also be vigilant regarding faithfulness to avoid repeating past mistakes (Jacovi & Goldberg, 2020). These new paradigms will present new arguments for why their method is faithful. As we are unfamiliar with these arguments, identifying their flaws is difficult, and it will be easy to get swayed by them.
Historically, a common tactic in post-hoc works was convincing visualization of explanations that aligned with our intuitions (Olah et al., 2017; Yosinski et al., 2015; Nguyen et al., 2016). However, such visualizations are empty arguments, as humans can’t know what a true explanation looks like (Geirhos et al., 2023). Likewise, intrinsic works have made seemingly strong theoretical arguments for why their methods were faithful, but these arguments failed to capture the whole model. Even the new learn-to-faithfully-explain paradigm have already shown sharp corners, where the explainer models unintentionally encode the prediction and are therefore unfaithful (Jethani et al., 2021).
To prevent false arguments, a sound start is to always have a specific and measurable definition of faithfulness, which works for all methods within a given explanation category (e.g., counterfactual or importance measures).
Finally, while these new paradigms are promising, it’s unlikely they will completely erase the current paradigms. We still teach both the particle and wave perspectives in physics. Most scientists don’t worry about whether there are true statements in math that cannot be proven.
Likewise, there will likely always be situations where intrinsic or post-hoc interpretability makes sense. For example, basic statistics and linear regressions can be framed as intrinsic interpretability. Hence, if a company or researchers decide to use a model because of its intrinsically explainable properties, then we should only praise them – as long as they also measure the faithfulness of the explanations.